OpenStudy (chris215):

Find all solutions in the interval [0, 2π). sin2 x + sin x = 0

OpenStudy (xapproachesinfinity):

hint sin2x=2sinxcosx

OpenStudy (xapproachesinfinity):

factor after that

OpenStudy (xapproachesinfinity):

\[\sin 2x+\sin x=0 \Longrightarrow 2\sin x\cos x+\sin x=0\] \[\sin x(2\cos x+1)=0 \] \[\sin x=0 ~~or~~ \cos x= \frac{-1}{2}\]

OpenStudy (xapproachesinfinity):

finish it up...